what are the properties of these
5+4=4+5
7(a-2)=7a-14
and
5(12)=5(14-2)

Answer by stanbon(75887)

5+4=4+5
commutative
----------------------
7(a-2)=7a-14
distributive
----------------------
and
5(12)=5(14-2)
identity
===========================
Cheers,
stan H.